# ALMAGAL VIII. Cataloging Hierarchical Mass Structure from Cores to Clumps across the Galactic Disk

### v1

@@ Abstract
Investigating the multi-scale fragmentation of dense clumps into compact cores is essential for understanding the processes that govern the initial distribution of mass in stellar clusters and how high-mass stars ($8~M_{\odot}$) form. We present a catalog of the hierarchical continuum structure from 904 clumps observed in the ALMAGAL program, a high resolution (0.150.80.15-0.8″) 1.38 mm Atacama Large Millimeter/submillimeter Array (ALMA) large program targeting dense clumps capable of high-mass star formation throughout the Galactic disk. We use astrodendro, a dendrogram-based algorithm, on a uniform linear resolution (2000 au) version of the data to extract 5160 continuum structures with effective radii spanning 80042000800-42000 au and estimated masses between $0.05-670~M_{\odot}$. With our large sample, we statistically examine the difference in clump properties for regions with varying levels of hierarchical complexity. We find that clumps exhibiting the richest hierarchical morphology have distributions with higher dust temperatures, surface densities, luminosity-to-mass (L/M) ratios, and most massive core (MMC) masses, indicating that these regions tend to be at later evolutionary stages. We find a positive correlation between the mass of cores from the ALMAGAL core catalog and the surface density of their surrounding structures identified in this work. However, this correlation is weaker for cores in more evolved clumps, where lower mass cores can be found at higher local surface densities. This could indicate that some cores accrete mass less efficiently from the intra-clump reservoir than others, despite the total available mass increasing over time, a scenario that is congruent with a clump-fed core accretion model.
@@
1 Introduction
Star formation occurs within a hierarchically structured interstellar medium composed of giant molecular clouds ($\sim$10–100 pc) that fragment into dense clumps ($\sim$ 1 pc) and then cores ($\sim$ 0.01–0.1 pc). These dense cores are the progenitors of individual stars or groups of stars (Lada & Lada, 2003). Approximately 50% of these stars form in densely populated (>>1000 stars) clusters containing at least one 10 M star (Adams, 2010). The stellar feedback from these high-mass stars is expected to have a large impact on the formation of stars in their immediate environment via gravitational interactions as well as mechanical and radiative feedback (e.g. Rosen & Krumholz, 2020).
There are competing theories on whether high-mass stars form directly from a massive ‘seed’ core, similar to their isolated, lower mass counterparts in a core-fed scenario (McKee & Tan, 2003; Tan & McKee, 2003; Krumholz et al., 2005, e.g.) or whether they dynamically accrete material from a larger scale mass reservoir in a clump-fed scenario (e.g. Zinnecker, 1982; Bonnell et al., 2001; Bonnell & Bate, 2006; Girichidis et al., 2011; Vázquez-Semadeni et al., 2019; Padoan et al., 2020; Anderson et al., 2021; Traficante et al., 2023). These dynamical accretion models differ in their expression. For example, in a competitive accretion model, cores that initially accrete mass from the intra-clump medium will continue to do so efficiently as their gravitational influence becomes more significant within the clump (Bonnell et al., 2001; Bonnell & Bate, 2006). However, in the fragmentation-induced starvation model, cores at the global minimum of the gravitational potential in a clump will accrete mass less efficiently as outer cores prevent some intra-clump gas from reaching the innermost cores (Peters et al., 2010; Girichidis et al., 2012).
@@ 1 Introduction
In this paper, we catalog the hierarchy of continuum structures observed in the ALMAGAL survey and investigate the multi-scale fragmentation of clumps across a range of evolutionary stages. It is organized as follows: in Section 2 we describe the ALMAGAL survey observations and the methods used to acquire our final data products. In Section 3 we outline the procedure used to catalog multi-scale continuum structures and the methods used for calculating their properties. We also define our method for classifying regions based on their relative level of hierarchical fragmentation. In Section 4 we report on the resulting continuum structure properties and show how the relative level of hierarchical structure seen in the ALMAGAL targets is correlated with its clump-scale properties. In Section 5 we discuss our results in the context of Galactic star formation, with a focus on how core mass accretion proceeds at different evolutionary stages. Finally, Section 6 is a brief summary of our conclusions.
Figure 1: Three ALMAGAL targets are selected as examples to depict typical morphology in the Isolated (top), Simple (middle), and Rich (bottom) clump categories. The method used to classify these targets is described in Section 3.2. The ID for each target is given at the top of each panel, and it corresponds to the “ALMAGAL ID” panel in Table 1. In the right panels, background grayscale fading into heat colorbar shows the ALMAGAL 1.3mm dust continuum emission. Purple contours indicate the branch structures identified by the dendrogram algorithm in each region, while white contours indicate the leaf structures. Cyan circles indicate the position of cores from the ALMAGAL core catalog (Coletta et al., 2025). The left panel in each figure shows the dendrogram structure produced from the images shown in the corresponding right panels using the methods described in Section 3.1. The size of the synthesized 2000 AU beam is shown in the bottom left corner of each image and a scale bar is included in the bottom right corner.
2 Data
@@
2.1 The ALMAGAL Survey Target Sample
The ALMAGAL target sample consists of 1013 dense clumps located throughout the Galactic disk. These regions are candidates for hosting high-mass star formation and were selected from 915 clumps identified in the Herschel Infrared Galactic Plane Survey (Hi-GAL, Molinari et al., 2010, 2016b; Elia et al., 2017, 2021), and 98 clumps from the Red MSX Survey (RMS, Lumsden et al., 2013). The initial proposed target selection criteria required that clumps have a mass >>500 $M_{\odot}$ for targets in the inner Galaxy and >>250 $M_{\odot}$ for those in the outer Galaxy. Observed targets were also required to have a surface density $\Sigma\geq$ 0.1 g cm-2, a threshold designed to include systems where high-mass stars are likely to form (Krumholz & McKee, 2008).
After a revision of the heliocentric distance estimates made after the ALMAGAL targets were observed, there are some targets that no longer fulfill these selection criteria. However, 90% of the targets still meet these initial selection requirements (Molinari et al., 2025). Target clumps span a wide range of heliocentric distance (0.5-10 kpc) and Galactocentric radii (3-14 kpc).
In Section 3.2, we evaluate the results from the continuum structure extraction as it relates to the properties of their host clumps, as they are reported in the revised and consolidated version of the ALMAGAL property table in Molinari et al. (2025) [^1] This table is available at the CDS via the following link: https://cdsarc.cds.unistra.fr/viz-bin/cat/J/A+A/696/A149. We note that there are 6 clumps for which an SED fit could not be used to make a dust temperature estimate, so they do not have reported masses or surface densities.
2.2 ALMAGAL Survey Observations
@@
2.3 Uniform Resolution ALMAGAL Continuum Maps
To ensure that our results are consistent for the entire ALMAGAL sample, we created a uniform linear resolution subset of the continuum images. First, we converted the native resolution images to the beam independent flux density unit of MJy/sr. Then, we calculated the circularized linear beam size for each image by taking the geometric mean of the major and minor axes of the beam. We then removed all targets with a corresponding linear spatial resolution >2000>2000 au, and smoothed the remaining 904 images with a 2D Gaussian kernel so that each source had a circularized beam with a linear size of 2000 au. We chose the 2000 au threshold to keep the majority of the full ALMAGAL sample ($\sim 90\%$), without reducing the spatial resolution of the data a substantial amount. In addition to this, we downsampled each smoothed image so that the 2000 au beam was sampled by 6 pixels. We note that only 900 of these targets have associated measured clump properties and consequently do not use the remaining 4 in the analysis presented in Section 4.2.
@@ 2.3 Uniform Resolution ALMAGAL Continuum Maps
(deg) (deg) (au) (Jy) (Jy) (MJy sr-1) (MJy sr-1)
0 AG022.7068+0.3897 277.806 -8.909 2348.9 2.56e-03 2.56e-03 5.74e+02 39.79 Leaf 0 Isolated
1 AG022.7068+0.3897 277.807 -8.909 2102.6 1.31e-03 1.31e-03 2.67e+02 39.79 Leaf 0 Isolated
2 AG022.7068+0.3897 277.807 -8.908 3826.5 4.18e-03 4.18e-03 2.47e+02 39.79 Leaf 0 Isolated
3 AG022.7384-0.2420 278.389 -9.176 1823.3 1.99e-03 1.99e-03 3.67e+02 50.03 Leaf 0 Simple
4 AG022.7384-0.2420 278.390 -9.174 10421.8 6.98e-02 6.98e-02 1.35e+03 50.03 Branch 0 Simple
5 AG022.7384-0.2420 278.390 -9.174 3300.5 1.59e-02 9.07e-03 1.35e+03 50.03 Leaf 2 Simple
6 AG022.7384-0.2420 278.390 -9.174 5489.4 3.88e-02 2.06e-02 1.35e+03 50.03 Branch 1 Simple
7 AG022.7384-0.2420 278.389 -9.174 4263.7 2.23e-02 1.10e-02 1.18e+03 50.03 Leaf 2 Simple
8 AG022.7384-0.2420 278.390 -9.174 2272.4 4.31e-03 1.09e-03 4.71e+02 50.03 Leaf 1 Simple
9 AG022.8624+0.4213 277.852 -8.757 5153.8 1.06e-02 1.06e-02 7.09e+02 48.70 Branch 0 Simple
2.4 The ALMAGAL Compact Core Catalog
@@
In total there are 6348 sources included in the ALMAGAL core catalog, with 5487 sources contained in our uniform linear resolution sample of 900 clumps. The physical size distribution is between 200 - 10000 au, with 90% of these fragments ranging from $\sim 800-3000$ au, with a median value of 1700 au. The calculated core masses span $0.002-345~M_{\odot}$, with a median value of 0.4 $M_{\odot}$. As described in Coletta et al. (2025), masses were calculated using the modified blackbody formula, assuming a dust opacity of 0.9 g cm-2 and a gas to dust ratio of 100. The core temperatures were estimated according to the $L/M$ ratio of its host clump (see Coletta et al. (2025) for details). We utilize this same prescription for calculating the mass of our dendrogram structures, as described in Section 3.3.
In this paper, we report the MMC mass from each clump using the core masses from the ALMAGAL compact core catalog.
2.5 Ancillary Data: ATLASGAL
@@
3 Methods
3.1 Dendrogram Algorithm
@@
We scale our inputs for these parameters using the rms noise ($\sigma_{\text{rms}}$) for each image so that $f_{\text{min}}=2\sigma_{\text{rms}}$, $f_{\text{peak}}=3\sigma_{\text{rms}}$, and $\delta=1\sigma_{\text{rms}}$. The $\sigma_{\text{rms}}$ noise values correspond to the ‘AGSTDREM’ values reported for the original, non-smoothed images from Sánchez-Monge et al. (2025), converted into units of MJy/sr. We take $n_{\text{pix}}$ to be equal to $0.5\times A_{\text{beam,pix}}$, where $A_{\text{beam,pix}}$ is the number of pixels contained within the Gaussian area of the synthesized beam for each image. We choose a factor of 0.5 instead of 1.0 because the area contained within the structure boundary at or above $f_{\text{min}}=2\sigma_{\text{rms}}$ may not exceed the full area of the beam, even if it is an otherwise significant detection. This can result in extracting structures that are slightly smaller than the beam, however this choice in $n_{\text{pix}}$ allows us to recover significant detections that would otherwise be removed. The first 10 rows of the catalog are shown in Table 1. In total we report 5160 structures in our catalog, 3470 of the structures are leaves and 1690 are branches. We present the dendrograms and corresponding structures for three different targets in Figure 1, where leaf and branch structures are shown in white and purple contours, respectively.
@@
3.2 Hierarchical Classification of the ALMAGAL Sample
We first consider the morphological complexity as a bulk property of each ALMAGAL target, or clump. As described in Section 3.1, each structure has a dendrogram ‘level’ that provides a relative description of how deeply nested the structure is in relation to its parent structure. For each source, we determine a maximum level, i.e. the level of the most deeply nested structure. If the maximum level of a region is relatively high, the region exhibits a more complex hierarchical structure.
@@ 3.2 Hierarchical Classification of the ALMAGAL Sample
We note that there are 48 Empty targets that have core detections from the ALMAGAL core catalog (Coletta et al., 2025) but no dendrogram detections. Additionally, there are 10 targets that have dendrogram detections, but no core detections, all of which are part of the Isolated sub-sample. We attribute this discrepancy to the difference in the data products and the structure identification algorithm used for this catalog and the ALMAGAL core catalog. For this catalog we use a spatially smoothed, common linear resolution version of the ALMAGAL data and a dendrogram-based extraction procedure, whereas the ALMAGAL core catalog is generated using the native resolution data products with the CuTEx algorithm.
@@
3.3 Calculation of Continuum Structure Properties
For each continuum structure in the catalog, we report its position, effective radius $R_{\text{eff}}$, integrated flux $F_{\nu}$, and peak intensity $I_{\text{peak}}$, as shown in Table 1. We calculate an effective projected radius from the total structure area: $R_{\text{eff}}\equiv(N_{\text{pix}}A_{\text{pix}}/\pi)^{1/2}$, where $N_{\text{pix}}$ is the total number of pixels in the structure, and $A_{\text{pix}}$ is the physical area of each pixel in au2. The integrated flux of the structure $F_{\nu}$ is obtained by summing the pixel values in the primary beam corrected image within the dendrogram structure and converting to Jy. For non-trunk structures we use the “clipping paradigm” (see Rosolowsky et al. 2008) to calculate the background subtracted integrated flux $F_{\nu\text{,bgs}}$ by subtracting the minimum intensity of the structure from each pixel and then summing over the entire structure.

### v2

# ALMAGAL VIII. Cataloging Hierarchical Continuum Structure from Cores to Clumps across the Galactic Disk

@@ Abstract
Investigating the multi-scale fragmentation of dense clumps into compact cores is essential for understanding the processes that govern the initial distribution of mass in stellar clusters and how high-mass stars ($8~M_{\odot}$) form. We present a catalog of the hierarchical continuum structure from 904 clumps observed in the ALMAGAL program, a high resolution (0.150.80.15-0.8″) 1.38 mm Atacama Large Millimeter/submillimeter Array (ALMA) large program targeting dense clumps capable of high-mass star formation throughout the Galactic disk. We use astrodendro, a dendrogram-based algorithm, on a uniform linear resolution (2000 au) version of the data to extract 5160 continuum structures with effective radii spanning 80042000800-42000 au and estimated masses between $0.05-670~M_{\odot}$. With our large sample, we statistically examine differences in clump properties for regions with varying levels of hierarchical complexity. We find that clumps exhibiting the richest hierarchical morphology have distributions with higher dust temperatures, surface densities, luminosity-to-mass (L/M) ratios, and most massive core (MMC) masses, indicating that these regions tend to be at later evolutionary stages. We find a positive correlation between the mass of cores from the ALMAGAL core catalog and the surface density of their surrounding structures identified in this work. However, this correlation is weaker for cores in more evolved clumps, where lower mass cores can be found at higher local surface densities. This could indicate that some cores accrete mass less efficiently from the intra-clump reservoir than others, despite the total available mass increasing over time, a scenario that is congruent with a clump-fed core accretion model.
@@
I Introduction
Star formation occurs within a hierarchically structured interstellar medium composed of giant molecular clouds ($\sim$10–100 pc) that fragment into dense clumps ($\sim$ 1 pc) and then cores ($\sim$ 0.01–0.1 pc). These dense cores are the progenitors of individual stars or groups of stars (Lada and Lada, 2003). Approximately 50% of these stars form in densely populated (>>1000 stars) clusters containing at least one 10 M star (Adams, 2010). The stellar feedback from these high-mass stars is expected to have a large impact on the formation of stars in their immediate environment via gravitational interactions as well as mechanical and radiative feedback (e.g. Rosen and Krumholz, 2020).
There are competing theories on whether high-mass stars form directly from a massive ‘seed’ core, similar to their isolated, lower mass counterparts in a core-fed scenario (McKee and Tan, 2003; Tan and McKee, 2003; Krumholz et al., 2005, e.g.) or whether they dynamically accrete material from a larger scale mass reservoir in a clump-fed scenario (e.g. Zinnecker, 1982; Bonnell et al., 2001; Bonnell and Bate, 2006; Girichidis et al., 2011; Vázquez-Semadeni et al., 2019; Padoan et al., 2020; Anderson et al., 2021; Traficante et al., 2023). These dynamical accretion models differ in their expression. For example, in a competitive accretion model, cores that initially accrete mass from the intra-clump medium will continue to do so efficiently as their gravitational influence becomes more significant within the clump (Bonnell et al., 2001; Bonnell and Bate, 2006). However, in the fragmentation-induced starvation model, cores at the global minimum of the gravitational potential in a clump will accrete mass less efficiently as outer cores prevent some intra-clump gas from reaching the innermost cores (Peters et al., 2010; Girichidis et al., 2012).
@@ I Introduction
In this paper, we catalog the hierarchy of continuum structures observed in the ALMAGAL survey and investigate the multi-scale fragmentation of clumps across a range of evolutionary stages. It is organized as follows: in Section II we describe the ALMAGAL survey observations and the methods used to acquire our final data products. In Section III we outline the procedure used to catalog multi-scale continuum structures and the methods used for calculating their properties. We also define our method for classifying regions based on their relative level of hierarchical fragmentation. In Section IV we report on the resulting continuum structure properties and show how the relative level of hierarchical structure seen in the ALMAGAL targets is correlated with its clump-scale properties. In Section V we discuss our results in the context of Galactic star formation, with a focus on how core mass accretion proceeds at different evolutionary stages. Finally, Section VI is a brief summary of our conclusions.
Figure 1: Three ALMAGAL targets are selected as examples to depict typical morphology in the Isolated (top), Simple (middle), and Rich (bottom) clump categories. The method used to classify these targets is described in Section III.2. The ID for each target is given at the top of each panel, and it corresponds to the “ALMAGAL ID” panel in Table 1. In the right panels, background grayscale fading into heat colorbar shows the ALMAGAL 1.3mm dust continuum emission. Purple contours indicate the branch structures identified by the dendrogram algorithm in each region, while white contours indicate the leaf structures. Cyan circles indicate the position of cores from the ALMAGAL core catalog (Coletta et al., 2025). The left panel in each figure shows the dendrogram structure produced from the images shown in the corresponding right panels using the methods described in Section III.1. The size of the synthesized 2000 AU beam is shown in the bottom left corner of each image and a scale bar is included in the bottom right corner.
II Data
@@
II.1 The ALMAGAL Survey Target Sample
The ALMAGAL target sample consists of 1013 dense clumps located throughout the Galactic disk. These regions are candidates for hosting high-mass star formation and were selected from 915 clumps identified in the Herschel Infrared Galactic Plane Survey (Hi-GAL, Molinari et al., 2010, 2016b; Elia et al., 2017, 2021), and 98 clumps from the Red MSX Survey (RMS, Lumsden et al., 2013). The initial proposed target selection criteria required that clumps have a mass >>500 $M_{\odot}$ for targets in the inner Galaxy and >>250 $M_{\odot}$ for those in the outer Galaxy. Observed targets were also required to have a surface density $\Sigma\geq$ 0.1 g cm-2, a threshold designed to include systems where high-mass stars are likely to form (Krumholz and McKee, 2008).
After a revision of the heliocentric distance estimates made after the ALMAGAL targets were observed, there are some targets that no longer fulfill these selection criteria. However, 90% of the targets still meet these initial selection requirements (Molinari et al., 2025). Target clumps span a wide range of heliocentric distances (0.5-10 kpc) and Galactocentric radii (3-14 kpc).
In Section III.2, we evaluate the results from the continuum structure extraction as it relates to the properties of their host clumps, as they are reported in the revised and consolidated version of the ALMAGAL property table in Molinari et al. (2025) [^1] This table is available at the CDS via the following link: https://cdsarc.cds.unistra.fr/viz-bin/cat/J/A+A/696/A149. We note that there are 6 clumps for which an SED fit could not be used to make a dust temperature estimate, so they do not have reported masses or surface densities.
II.2 ALMAGAL Survey Observations
@@
II.3 Uniform Resolution ALMAGAL Continuum Maps
To ensure that our results are consistent for the entire ALMAGAL sample, we created a uniform linear resolution subset of the continuum images. First, we converted the native resolution images to the beam independent flux density unit of MJy/sr. Then, we calculated the circularized linear beam size for each image by taking the geometric mean of the major and minor axes of the beam. We then removed all targets with a corresponding linear spatial resolution >2000>2000 au, and smoothed the remaining 904 images with a 2D Gaussian kernel so that each source had a circularized beam with a linear size of 2000 au. We chose the 2000 au threshold to keep the majority of the full ALMAGAL sample ($\sim 90\%$), without reducing the spatial resolution of the data by a substantial amount. Additionally, we downsampled each smoothed image so that the 2000 au beam was sampled by 6 pixels. We note that only 900 of these targets have associated measured clump properties and consequently do not use the remaining 4 in the analysis presented in Section IV.2.
@@ II.3 Uniform Resolution ALMAGAL Continuum Maps
(deg) (deg) (au) (mJy) (mJy) (MJy sr-1) (MJy sr-1)
0 AG022.7068+0.3897 277.806 -8.909 2349 2.56 2.56 5.74e+02 39.79 Leaf 0 isolated
1 AG022.7068+0.3897 277.807 -8.909 2103 1.31 1.31 2.67e+02 39.79 Leaf 0 isolated
2 AG022.7068+0.3897 277.807 -8.908 3827 4.18 4.18 2.47e+02 39.79 Leaf 0 isolated
3 AG022.7384-0.2420 278.389 -9.176 1823 1.99 1.99 3.67e+02 50.03 Leaf 0 simple
4 AG022.7384-0.2420 278.390 -9.174 10422 69.76 69.76 1.35e+03 50.03 Branch 0 simple
5 AG022.7384-0.2420 278.390 -9.174 3300 15.85 9.07 1.35e+03 50.03 Leaf 2 simple
6 AG022.7384-0.2420 278.390 -9.174 5489 38.82 20.63 1.35e+03 50.03 Branch 1 simple
7 AG022.7384-0.2420 278.389 -9.174 4264 22.31 11.02 1.18e+03 50.03 Leaf 2 simple
8 AG022.7384-0.2420 278.390 -9.174 2272 4.31 1.09 4.71e+02 50.03 Leaf 1 simple
9 AG022.8624+0.4213 277.852 -8.757 5154 10.63 10.63 7.09e+02 48.70 Branch 0 simple
II.4 The ALMAGAL Compact Core Catalog
@@
In total there are 6348 sources included in the ALMAGAL core catalog, with 5487 sources contained in our uniform linear resolution sample of 900 clumps. The physical size distribution is between 200 - 10000 au, with 90% of cores ranging from $\sim 800-3000$ au, with a median value of 1700 au. The calculated core masses span $0.002-345~M_{\odot}$, with a median value of 0.4 $M_{\odot}$. As described in Coletta et al. (2025), masses were calculated using the modified blackbody formula, assuming a dust opacity of 0.9 g cm-2 and a gas to dust ratio of 100. The core temperatures were estimated according to the $L/M$ ratio of its host clump (see Coletta et al. (2025) for details). We utilize this same prescription for calculating the mass of our dendrogram structures, as described in Section III.3.
In this paper, we report the MMC mass from each clump using the core masses from the ALMAGAL compact core catalog. Additionally, we note that we use the terminology of “core” to specifically refer to the cores from the ALMAGAL core catalog, and not objects from the hierarchical structure catalog presented here.
II.5 Ancillary Data: ATLASGAL
@@
III Methods
III.1 Dendrogram Algorithm
@@
We scale our inputs for these parameters using the rms noise ($\sigma_{\text{rms}}$) for each image so that $f_{\text{min}}=2\sigma_{\text{rms}}$, $f_{\text{peak}}=3\sigma_{\text{rms}}$, and $\delta=3\sigma_{\text{rms}}$. The $\sigma_{\text{rms}}$ noise values correspond to the ‘AGSTDREM’ values reported for the original, non-smoothed images from Sánchez-Monge et al. (2025), converted into units of MJy/sr. We take $n_{\text{pix}}$ to be equal to $0.5\times A_{\text{beam,pix}}$, where $A_{\text{beam,pix}}$ is the number of pixels contained within the Gaussian area of the synthesized beam for each image. We choose a factor of 0.5 instead of 1.0 because the area contained within the structure boundary at or above $f_{\text{min}}=2\sigma_{\text{rms}}$ may not exceed the full area of the beam, even if it is an otherwise significant detection. This can result in extracting structures that are slightly smaller than the beam, however this choice in $n_{\text{pix}}$ allows us to recover significant detections that would otherwise be removed. The first 10 rows of the catalog are shown in Table 1. In total we report 5160 structures in our catalog, 3470 of the structures are leaves and 1690 are branches. We present the dendrograms and corresponding structures for three different targets in Figure 1, where leaf and branch structures are shown in white and purple contours, respectively.
@@
All 391 280 229 900
Near 213 112 119 444
Far 178 168 110 456
III.2 Hierarchical Classification of the ALMAGAL Sample
We first consider the morphological complexity as a bulk property of each ALMAGAL target, or clump. As described in Section III.1, each structure has a dendrogram ‘level’ that provides a relative description of how deeply nested the structure is in relation to its parent structure. For each source, we determine a maximum level, i.e. the level of the most deeply nested structure. If the maximum level of a region is relatively high, the region exhibits a more complex hierarchical structure.
@@ III.2 Hierarchical Classification of the ALMAGAL Sample
We note that there are 48 Empty targets that have core detections from the ALMAGAL core catalog (Coletta et al., 2025) but no dendrogram detections. Additionally, there are 10 targets that have dendrogram detections, but no core detections, all of which are part of the Isolated sub-sample. We attribute this discrepancy to the difference in the data products and the structure identification algorithm used for this catalog and the ALMAGAL core catalog. For this catalog we use a spatially smoothed, common linear resolution version of the ALMAGAL data and a dendrogram-based extraction procedure, whereas the ALMAGAL core catalog is generated using the native resolution data products with the CuTEx algorithm. Additionally, for the hierarchical structure catalog, a structure is required to have a peak intensity of $3\sigma_{\text{rms}}$ and a minimum intensity threshold of $2\sigma_{\text{rms}}$ at the structure boundary. For the ALMAGAL core catalog, core detections must have a peak intensity of at least $5\sigma_{\text{rms}}$ and there is no minimum intensity threshold imposed.
@@
III.3 Calculation of Continuum Structure Properties
For each continuum structure in the catalog, we report its position, effective radius $R_{\text{eff}}$, integrated flux $F_{\nu}$, and peak intensity $I_{\text{peak}}$, as shown in Table 1. The coordinates of the structure correspond to the average position of the continuum structure, as measured within the dendrogram contour. We calculate an effective projected radius from the total structure area: $R_{\text{eff}}\equiv(N_{\text{pix}}A_{\text{pix}}/\pi)^{1/2}$, where $N_{\text{pix}}$ is the total number of pixels in the structure, and $A_{\text{pix}}$ is the physical area of each pixel in au2. The integrated flux of the structure $F_{\nu}$ is obtained by summing the pixel values in the primary beam corrected image within the dendrogram structure and converting to Jy. For non-trunk structures we use the “clipping paradigm” (see Rosolowsky et al. 2008) to calculate the background subtracted integrated flux $F_{\nu\text{,bgs}}$ by subtracting the minimum intensity of the structure from each pixel and then summing over the entire structure.